### v2

@@ Abstract
Logs produced by extensive software systems are integral to monitoring system behaviors. Advanced log analysis facilitates the detection, alerting, and diagnosis of system faults. Log parsing, which entails transforming raw log messages into structured templates, constitutes a critical phase in the automation of log analytics. Existing log parsers fail to identify the correct templates due to reliance on human-made rules. Besides, These methods focus on statistical features while ignoring semantic information in log messages. To address these challenges, we introduce a cutting-edge Log parsing framework with Entropy sampling and Chain-of-Thought Merging (Lemur). Specifically, to discard the tedious manual rules. We propose a novel sampling method inspired by information entropy, which efficiently clusters typical logs. Furthermore, to enhance the merging of log templates, we design a chain-of-thought method for large language models (LLMs). LLMs exhibit exceptional semantic comprehension, deftly distinguishing between parameters and invariant tokens. We have conducted experiments on large-scale public datasets. Extensive evaluation demonstrates that Lemur achieves the state-of-the-art performance and impressive efficiency. [^1] https://github.com/knediny/lemur.
@@ 1 Introduction
Logs serve as a critical information source for system monitoring, offering key insights into system behavior, a fact well-documented in existing literature He et al. (2021). Their advantage over other types of data lies in their rich informational content and relative ease of interpretation. Through log analysis, several important downstream tasks can be effectively addressed. Which includes anomaly detection Du et al. (2017); Nandi et al. (2016), fault diagnosis He et al. (2018); Zou et al. (2016), and root cause analysis Tak et al. (2016). Log parsing, a crucial initial step in log analysis He et al. (2016), separates log messages into two elements: 1) Log Templates. The constant, unchanging parts in logging statements, and 2) Log Variables. The dynamic, changeable details in different executions. In Figure 1, the logging statement logger.info(’Waited {waited_time} seconds for {process} to be killed.’) can yield various messages like Wait 26 seconds for Thread-20 to be killed. Waited <*><*>< * > seconds for <*><*>< * > to be killed is the log template, and the changing data like 26 and Thread-20 are the log variables.
Figure 1: An example of log parsing process.
@@
In the realm of system maintenance, where source code accessibility is often restricted, different parsers He et al. (2017); Fu et al. (2022); Le and Zhang (2023); Yu et al. (2023); Jiang et al. (2023) have been developed to autonomously differentiate between templates and parameters within log messages. Existing log parsers can be categorized into two groups: syntax-based and semantic-based. Syntax-based log parsers He et al. (2017); Du and Li (2016); Makanju et al. (2009); Vaarandi (2003); Vaarandi and Pihelgas (2015) utilize specific features or heuristics (e.g., word frequency) to extract the constant parts of log messages as templates. Conversely, semantic parsers Liu et al. (2022); Le and Zhang (2023) leverage advanced deep learning frameworks to assimilate semantics and system-specific patterns from annotated log data, thereby facilitating the parsing of new log entries. Besides, recent works Jiang et al. (2023); Guo et al. (2023) leverage large language models (LLMs) OpenAI (2023) for log parsing, which also utilizes the powerful semantic understanding of LLMs.
@@
Nonetheless, significant challenges persist. Syntax-based methodologies are heavily dependent on meticulously crafted rules. Their performance significantly diminishes with the exponential increase in the volume and intricacy of log data. Furthermore, these techniques often overlook semantic variances among logs. For instance, log entries such as ’success to open file/var/log/system’ and ’success to close file/var/log/system’ display syntactic resemblance, potentially leading to their aggregation by data-driven approaches, which could adversely impact the efficacy of anomaly detection tasks. Semantic-based methods, reliant on pre-labeled logs for training, falter in understanding semantics when applied to unfamiliar domains. LLM-based parsers Liu et al. (2023); Le and Zhang (2023); Jiang et al. (2023) are impractical due to high inference time (such as GPUs) and network latency Jiao et al. (2023). Besides, LLMs generate unstable results because of the illusion problem.
@@
To tackle these challenges, we propose Lemur, a cutting-edge Log parsing framework with Entropy sampling and Chain-of-Thought Merging (Lemur). Lemur brings together the strengths of the syntax-based and semantic-based methods, which consists of three key components: Information Entropy Clustering, Template Generation and Chain-of-Thought Merging. Specifically, inspired by information theory Gray (2011), we recognize that different logs encapsulate varying quantities of information, while logs of a similar nature contain comparable amounts. Consequently, we have developed a novel sampling method grounded in the principles of information entropy, which efficiently clusters characteristic logs. Then in template generation, We determine the variables in the log based on the information entropy of the token at the same location. In the end, motivated by Chain-of-Thought Wei et al. (2022a), we design a three-hop Chain-of-Thought (infer structure, infer semantics and infer solution) for merging templates.
@@
We have conducted a comprehensive evaluation on public large-scale log datasets of LogHub He et al. (2020) with seven state-of-the-art log parsing approaches. The results show that Lemur achieves the highest performance than other baselines for the F1 score of grouping and template accuracy. Generally, the main contributions of this work are listed as follows:
@@
  • We introduce a novel sampling method based on information entropy for efficient log clustering. Besides, utilizing LLMs, we can accurately merge templates based on semantics than synax-based methods.
  • Extensive experiments are conducted on public benchmark to demonstrate the effectiveness of our Lemur. The results show that Lemur outperforms other state-of-the-art methods.
2 Related Work
2.1 Log Parser
In the evolving field of automatic log analysis, crucial for distributed systems and cloud computing, significant progress has been made in log parsing techniques, categorized into frequent pattern mining, clustering, and heuristics rules. Frequent pattern mining is exemplified by SLCT Vaarandi (2003) which groups logs based on token frequency, and LogCluster Vaarandi and Pihelgas (2015) that removes positional constraints in log grouping.. Clustering approaches include LogMine Hamooni et al. (2016) with its multi-layered clustering system, LKEFu et al. (2009) using edit distance and position weighing, SHISO Mizutani (2013) improving efficiency through hierarchical clustering, LenMa Shima (2016) employing token length vectors for clustering, and LPV Xiao et al. (2020) which uses semantic vectors from word2vec. In the heuristics rules category, IPLOM Makanju et al. (2009) partitions logs by length and token position, Spell Du and Li (2016) approaches parsing as a longest common sequential problem, Drain He et al. (2017) groups logs by length and prefixes for template updates, and Prefix-Graph Chu et al. (2021) merges prefix trees into graphs for template generation. Recent advancements have introduced deep learning-based algorithms like Nulog Nedelkoski et al. (2021) with a transformer encoder layer and masked language model, Uniparser Liu et al. (2022) comprising multiple modules for comparative learning, and LogAP Rand and Miranskyy (2021) utilizing machine translation for parsing. However, these deep learning methods face challenges in efficiency and high operational costs due to GPU requirements.
2.2 Large Language Model
Language Models.
Language, a unique human ability developing from early childhood, necessitates advanced artificial intelligence algorithms for machines to comprehend and utilize it. Language modeling using self-supervised learning and large-scale data, particularly through pre-training large Transformer encoders and decoders Vaswani et al. (2017); Chi et al. (2020), significantly enhances various natural language processing tasks. Specifically, pre-training a Transformer decoder OpenAI (2023) aids in unconditional text generation. Performance improvements in diverse tasks have been linked to the enlargement of Pre-training Language Models (PLMs) by increasing model or data size, following established scaling principles. This has led to the creation of increasingly larger PLMs, such as GPT-3 with 175 billion parameters and PaLM with 540 billion Anil et al. (2023), guided by the scaling laws of large language models Aghajanyan et al. (2023).
Despite their similar architectures and pre-training tasks, larger PLMs, such as GPT-4 OpenAI (2023), exhibit unique behaviors and emergent abilities, excelling in complex tasks. A prime example is ChatGPT, adapting GPT-series LLMs for engaging dialogues, showcasing advanced conversational skills. Fine-tuning LLMs on various datasets Chung et al. (2022); Wei et al. (2022b) yields promising results, using human or LLM-created prompts for instruction tuning and refining generations. Chain-of-thought prompting Wei et al. (2022b), where models explain their reasoning for complex problems to improve answer accuracy, and RLHF Ouyang et al. (2022), a strategy for fine-tuning LLMs, significantly enhance their performance.
@@ 3 Methodology
In this section, we provide a detailed and thorough overview of Lemur, a framework specifically engineered to significantly improve the accuracy and flexibility of log message categorization. Illustrated in Figure 2, Lemur is meticulously structured into three primary components: Information Entropy Clustering, Template Generation, and Chain-of-Thought Merging. The discussion commences with an in-depth look at the data preprocessing stage, followed by an elaborate exposition on the intricacies and functionalities of each component.
Figure 2: Overview of our log parsing framework
3.1 Data Preprocessing
Effective log data preprocessing is an unavoidable component for log parsing and vigilant monitoring of systems. This process involves a series of orchestrated steps that transform raw logs into meticulously semi-structured data. In this stage, we need to extract vital content, replace identifiable variables with uniform tokens, and remove useless characters.
Firstly, we construct a regular expression tailored to the format of the original log entries to distill the main messages. For example, HDFS dataset provided by Loghub adheres to a specific template: <Date> <Time> <Pid> <Level> <Component> : <Content>. The <Content> part warrants in-depth investigation while the rest serve as contextual scaffolding. Then identifiable variables including IP address or URL need to be replaced with generic token <*>. After that, we remove the quotation marks at the beginning and the end of the content.
@@ 3.2 Information Entropy Clustering
After pre-processing, we gain a set of logs $L$ = $\{L_{1},L_{2},\ldots,L_{N}\}$. Then we determine the set of all unique lengths from these logs, denoted as $l$ = $\{l_{1},l_{2},\ldots,l_{M}\}$, where $M$ is the number of unique lengths. Based on $l$, we create $M$ buckets, denoted as $B$ = $\{B_{1},B_{2},\ldots,B_{M}\}$, where bucket $B_{j}$ is intended to store logs of length $l_{j}$. The process can be described by a mapping function $f$:
$f=\{L_{1},L_{2},\ldots,L_{N}\}\rightarrow\{B_{1},B_{2},\ldots,B_{M}\}$ (1)
Thus, each log is allocated to a distinct bucket based on its length. This methodical sorting is crucial for maintaining analytical precision as it minimizes intra-bucket variability, which can otherwise lead to decreased clustering performance.
Entropy Sampling
In order to find possible log templates from log token list buckets with the same token length, a natural idea is to cluster them. Different from common clustering methods that randomly select and then gradually change, inspired by information theory Gray (2011), we have developed a novel sampling method grounded in the principles of information entropy, which efficiently clusters characteristic logs. For each bucket $B_{j}\in B$, we select distinct log samples based on entropy theory Gray (2011) as the clustering centers. Suppose we select $k$ samples total. For each log $x\in B_{j}$, we need to compute its entropy $E(x)$ using the Shannon entropy formula Gray (2011):
$E(x)=-\sum_{i}p(x_{i})\log p(x_{i})$ (2)
where $x_{i}$ is the token in log $x$. Logs are sorted in descending order based on their $E(x)$ and then divided into $n$ layers $\{Lay_{1},Lay_{2},\ldots,Lay_{n}\}$. For each layer $Lay_{i}$, the process iterates through the logs. A log $x$ is added to the selected set $S$ if its first token $T(x)$ has not been selected before, or if the number of selected tokens $R$ is at least $k$. The process stops if the size of samples set $S$ reaches $k$:
$\begin{cases}x\in S,&\text{if }(T(x)\notin S_{T}\lor R\geq k)\land|S_{T}|\geq k,\\ \text{Stop},&\text{if }|S|\geq k.\end{cases}$ (3)
where $S_{T}$ is the set of selected first tokens of $x$. $R$ denotes the number of selected tokens. After we gain $S$ set, we merge the $K$ log centers based on Jaccard similarity in Algorithm 1. Let $J_{T}$ as the threshold of the Jaccard similarity, if If the similarity between the two centers ($L_{1},L_{2}$) is greater than $J_{T}$, we remove $L_{2}$. After merging, we get $k^{{}^{\prime}}$ samples as new cluster centers.
Algorithm 1 Jaccard Similarity in Two Log Token Lists
Input: $L_{1},L_{2}$ (Logs), $J_{T}$ (threshold)
Output: $similarity$
1:  $M1\leftarrow\text{SET}(L_{1})$
2:  $M2\leftarrow\text{SET}(L_{2})$
3:  $similarity\leftarrow\frac{|M1\cap M2|}{|M1\cup M2|}$
4:  if $similarity>J_{T}$ then
5:     Remove $L_{2}$
6:  end if
7:  return $similarity$
@@
Clustering in modern systems is a critical method for managing large-scale log data, simplifying analysis by grouping similar logs. Clustering not only reduces complexity and time involved in direct analysis but also facilitates the discovery of hidden patterns related to system performance and security. Additionally, it aids in template generation, distinguishing between static and dynamic log components, thereby streamlining data interpretation and monitoring.
For each bucket $B_{j}\in B$, we achieve S as clustering centers. Suppose $x_{c}$ be the number of cluster centers, each denoted as $C_{i}\in S$ $(i=1,2,\ldots,x_{c})$. $x$ be the number of logs in $B_{j}$, each denoted as $L_{j}$ $(j=1,2,\ldots,x)$. For each log entry $L_{j}$ and each cluster center $C_{i}$, the distance $D(L_{j},C_{i})$ is defined as:
$D(L_{j},C_{i})=\sum_{k=1}^{n}\mathbb{I}(L_{jk}\neq C_{ik})$ (4)
Each log entry $L_{j}$ is assigned to the $C_{i}$ with the minimum distance:
[…]
@@ 3.3 Template Generation
In template generation, our goal is to identify the variables in logs. First, we utilize the Longest Common Subsequence (LCS) among all logs, which may either be non-empty or empty. Subsequently, For each log, we compare it with the LCS and record the location of the token that is not the same. Following this, an exhaustive analysis is conducted over the positions of all divergent tokens, during which the information entropy and the list of tokens at these positions are calculated for all logs. This methodical approach facilitates a nuanced understanding of the variances within the log data, providing critical insights for comprehensive log analysis. The process of finding variables are in the following:
$\text{VPD}=\begin{cases}\text{Yes},&\text{if }-\sum_{t\in T}\frac{f(t)}{N}\log_{2}\frac{f(t)}{N}>\theta\\ \text{No},&\text{if }-\sum_{t\in T}\frac{f(t)}{N}\log_{2}\frac{f(t)}{N}\leq\theta\end{cases}$ (6)
where $VPD$ donates variation point determination, The calculation of entropy $H$ is based on the distribution of tokens occurring at a specific position. $T$ is the set of tokens at a specific position. $f(t)$ is the frequency of token $t$ in all logs. $N$ is the total number of logs. $\theta$ is the threshold value. The decision for variables is $Yes$ if the calculated entropy exceeds the threshold $\theta$, otherwise $No$.
@@ 3.4 Three-S-hop Chain-of-Thought Merging
In the subsequent phase of log message analysis, following the segregation of messages into distinct clusters and the delineation of unique log templates, a notable challenge emerges in the reconciliation of templates of varying lengths that, despite their differences, may semantically represent identical events. This issue primarily stems from the limitation of length-based classification methods, which often fail to acknowledge that tokens of semantic equivalence could correspond to a singular template entity. To address this challenge effectively, we incorporate the sophisticated semantic parsing capabilities inherent in large language models (LLMs). These models are particularly adept at interpreting textual data within its contextual framework, enhancing the accuracy and efficacy of template reconciliation in log analysis.
In Figure 3, our Chain-of-Thought Merging consists of three-hop inferrring: structure, semantics and solution. Specially, we initially pinpoint pairs of log templates that, although distinct in length, are likely to encapsulate the same event. Subsequently, for each identified pair, we meticulously craft a series of prompts designed to encapsulate the nuances of their individual semantic contexts. These prompts, exemplified in Figure 3, are then fed into a large language model. It triggers the merging of the templates into a unified template when the templates obtain a substantial degree of similarity. This newly formed template adeptly encapsulates the core of the event, transcending the variations in token count. This process of merging is methodically applied across all identified template pairings, culminating in a refined and consolidated set of log templates. These templates adeptly span the spectrum of logged events, enhancing the overall accuracy and coherence of the log parsing framework.
Figure 3: Three-S-Hop Chain-of-Thought Merging
4 Experiment
4.1 Implement Details
Datasets
Experiments are conducted on the most widely-used benchmark datasets published in LogHubHe et al. (2020). More details are available in Table 1.
Dataset Templates Avg. Len. Various Lengths Proportion Messages
@@ 4.1 Implement Details
Figure 5: Boxplot of group accuracy on 16 benchmark datasets.
Implementation and Configuration
We implement Lemur based on Python 3.10, Apple M3 chip with 8-core CPU, 10-core GPU and 16GB RAM, macOS Sonoma(14.2.1). In experiments, $D$ is the tokens for word-splitting, $k$ denotes the number of the $S$ in each bucket, the Jaccard similarity threshold for within-bucket merging $J_{T}$ and the entropy threshold for token identification $\theta$ are shown in Table 2.
@@
4.2 Baselines and Metrics
Baselines
As for baselines, We choose DrainHe et al. (2017), SpellDu and Li (2016), IPLOMMakanju et al. (2009), ULPSedki et al. (2022), BrainYu et al. (2023), LogPPTLe and Zhang (2023) and LLMParserJiang et al. (2023) as our baselines.
F1 score of Grouping Accuracy (FGA).
Grouping Accuracy (GA).
Figure 6: Execution time for various datasets.
Figure 7: Execution time for various dataset size.
Execution time.
4.3 Main Results
In Table 3, Lemur, an unsupervised LLM-related model, exhibits remarkable effectiveness and performance in the field of log parsing. When compared to Brain, the current state-of-the-art in unsupervised models, Lemur demonstrates superior or comparable results across various datasets, underscoring its efficacy in unsupervised learning without the need for annotated data. Moreover, even when juxtaposed with supervised models like LogPPT and LLMParser, Lemur shows equal or better performance in most datasets, based on F-measure and GA metrics. This is particularly noteworthy as it highlights the capability of Lemur to match or exceed the performance levels of supervised models, despite the absence of explicit label guidance.

### v3

@@ Abstract
Logs produced by extensive software systems are integral to monitoring system behaviors. Advanced log analysis facilitates the detection, alerting, and diagnosis of system faults. Log parsing, which entails transforming raw log messages into structured templates, constitutes a critical phase in the automation of log analytics. Existing log parsers fail to identify the correct templates due to reliance on human-made rules. Besides, these methods focus on statistical features while ignoring semantic information in log messages. To address these challenges, we introduce a cutting-edge Log parsing framework with Entropy sampling and chain-of-thought Merging (Lemur). Specifically, to discard the tedious manual rules, we propose a novel sampling method inspired by information entropy, which efficiently clusters typical logs. Furthermore, to enhance the merging of log templates, we design a chain-of-thought method for large language models (LLMs). LLMs exhibit exceptional semantic comprehension and deftly distinguish between parameters and invariant tokens. We have conducted experiments on large-scale public datasets. Extensive evaluation demonstrates that Lemur achieves state-of-the-art performance and impressive efficiency. The Code is available at https://github.com/zwpride/lemur.
Keywords:
Log Parsing Large Language Models Information Entropy.
@@ 1 Introduction
Logs serve as a critical information source for system monitoring, offering key insights into system behavior, a fact well-documented in existing literature [12]. Their advantage over other types of data lies in their rich informational content and relative ease of interpretation. Through log analysis, several important downstream tasks can be effectively addressed, which include anomaly detection [5, 23], fault diagnosis [13, 39], and root cause analysis [31]. Log parsing, a crucial initial step in log analysis [10], separates log messages into two parts: 1) Log Templates. The constant, unchanging parts in logging statements, and 2) Log Variables. The dynamic, changeable details in different executions. In Figure 1, the logging statement logger.info(’Wait {wait_time} seconds for {process} to be killed.’) can yield various messages like Wait 26 seconds for Thread-20 to be killed. Wait <><*>< ∗ > seconds for <><*>< ∗ > to be killed is the log template, and the changing data like 26 and Thread-20 are the log variables.
Figure 1: An example of log parsing. Logging Statement cannot be accessed in most cases. Log Message is parsed into a static Template part containing fixed keywords and a Variable part that will vary between different log messages with the same template.
@@
In the field of system maintenance, where source code accessibility is often restricted, different parsers [11, 7, 18, 38, 15] have been developed to differentiate between templates and variables within log messages autonomously. Existing log parsers can be categorized into two groups: syntax-based and semantic-based. Syntax-based log parsers [11, 4, 21, 32, 33] utilize specific features or heuristics (e.g., word frequency) to extract the constant parts of log messages as templates. Conversely, semantic-based parsers [20, 18] leverages advanced deep learning frameworks to assimilate semantics and system-specific patterns from annotated log data, thereby facilitating the parsing of new logs. Besides, recent works [15] leverage large language models (LLMs) [25] for log parsing, which also utilizes the powerful semantic understanding of LLMs.
@@
However, syntax-based methodologies are heavily dependent on meticulously crafted rules. Their performance significantly diminishes with the exponential increase in the volume and intricacy of log data. Furthermore, these techniques often overlook semantic variances among logs. For instance, logs such as success to open file/var/log/system and success to close file/var/log/system display syntactic resemblance, potentially leading to their aggregation by data-driven approaches, which could adversely impact the efficacy of anomaly detection tasks. Semantic-based methods, reliant on pre-labeled logs for training, falter in understanding semantics when applied to unfamiliar domains. LLM-based parsers [19, 18, 15] have lower availability due to high inference time (such as GPUs) and network latency [16]. Besides, LLMs generate unstable results because of the hallucination problem.
@@
To tackle these challenges, we propose Lemur, a cutting-edge Log parsing framework with Entropy sampling and Chain-of-Thought Merging (Lemur). Lemur brings together the strengths of the syntax-based and semantic-based methods, which consist of three key components: Information Entropy Clustering, Template Generation, and Chain-of-Thought Merging. Specifically, inspired by information theory [8], we recognize that different logs encapsulate varying quantities of information, while logs of a similar nature contain comparable amounts. Consequently, we have developed a novel sampling method based on information entropy principles that efficiently clusters characteristic logs by dividing large-scale data into multiple clusters and utilizing efficient sampling and clustering algorithms within those clusters, thereby ensuring that LEMUR remains robust and high-performance in large-scale log scenarios. Then in template generation, we determine the variables and the template in the log based on the information entropy of the token at the same location. In the end, motivated by Chain-of-Thought [35], we design a three-hop Chain-of-Thought (infer structure, infer semantics, and infer solution) for merging templates.
Figure 2: Overview of our log parsing framework.
@@
We have conducted a comprehensive evaluation on public large-scale log datasets of LogHub [14] with seven state-of-the-art log parsing approaches. The results show that Lemur achieves the highest performance than other baselines for the F1 score of grouping and template accuracy. Generally, the main contributions of this work are listed as follows:
@@
  • We introduce a novel sampling method based on information entropy for efficient log clustering. Besides, by utilizing LLMs, we can accurately merge templates based on semantics rather than syntax-based methods.
  • Extensive experiments are conducted on public benchmarks to demonstrate the effectiveness of our Lemur. The results show that Lemur outperforms other state-of-the-art methods.
@@ 2 Methodology
In this section, we overview Lemur, a framework designed to enhance log message categorization. Figure 2 shows Lemur, which consists of three main components: Information Entropy Clustering, Template Generation, and Chain-of-Thought Merging.
@@ 2.1 Information Entropy Clustering
In this section, logs are categorized into buckets based on their length. Some representative logs are selected as the cluster centers within each bucket, and other logs are clustered around these centers.
@@
Logs are assigned to buckets $B_{j}$ by a mapping function $f$ based on length: $f:L\rightarrow B$, with $L$ representing logs and $B$ representing buckets. For each log $L_{i}\in L$, we compute its length $len(L_{i})$ after tokenization. The log is then assigned to bucket $B_{j}\in B$, where $j$ matches the index of the log’s length in the set of unique lengths $l$. This can be written as $f(L_{i})=B_{j}$ where $l_{j}=len(L_{i})$. This method ensures logs of the same length are grouped, reducing variability and aiding subsequent analysis.
Entropy-based Sampling
To identify potential log templates from buckets of logs with identical token lengths, we propose a clustering method inspired by information theory [8], specifically employing Shannon entropy. Unlike conventional clustering which uses random seeds and iterative refinement, our approach selects logs based on their information content, which is ideal for log analysis due to the variability in log messages.
Each log $x\in B_{j}$ is evaluated by its entropy $E(x)=-\sum_{i}p(x_{i})\log p(x_{i})$, where $x_{i}$ is the token and $p(x_{i})$ its probability, gauging its information content. Logs are then ranked by descending entropy into layers ${Lay_{1},\ldots,Lay_{n}}$ to prioritize those with rich information.
We select clustering centers from these layers, starting with the highest entropy logs and picking those with either new first tokens or entropy above a threshold. This process repeats until we’ve chosen $k$ centers or the remaining logs no longer offer new information. Our stopping criterion ensures we gather diverse and informative logs while avoiding redundancy.
Refinement with Jaccard Similarity
After obtaining the initial set of $k$ samples, we further refine the selection by merging similar log centers based on their Jaccard similarity. This merging process helps eliminate redundancy while preserving the diversity of log samples. We use the Jaccard similarity to measure the similarity between two logs. The Jaccard similarity between two logs $L_{1}$ and $L_{2}$ is calculated as $similarity=\frac{|L_{1}\cap L_{2}|}{|L_{1}\cup L_{2}|}$. Let $J_{T}$ be the threshold for Jaccard similarity; if the similarity between two centers ($L_{1}$, $L_{2}$) exceeds $J_{T}$, we remove $L_{2}$. This refinement ensures a diverse and representative set of $k^{\prime}$ log samples and optimizes the balance between coverage and conciseness in the log samples.
@@
In the clustering process for logs within a bucket $B_{j}\in B$, we distinguish between the set of representative logs, $S$, which serve as cluster centers, and the set of remaining logs, $O$, which are to be clustered around these centers.
Let $S=\{s_{1},s_{2},\ldots,s_{k}\}$ represent the cluster centers and $O=\{o_{1},o_{2},\ldots,o_{n}\}$ denote the logs to be clustered. The objective is to assign each log $o_{i}\in O$ to a cluster center $s_{j}\in S$ such that the edit distance between them, $d(o_{i},s_{j})$, is minimized. The edit distance measures the similarity between two logs in terms of the changes required to convert one into the other at the token level. The edit distance $d(o_{i},s_{j})$ between a log $o_{i}$ and a cluster center $s_{j}$ is defined as:
$d(o_{i},s_{j})=\min_{s_{j}\in S}\sum_{t=1}^{T}\delta(o_{i,t},s_{j,t})$
where $\delta(o_{i,t},s_{j,t})$ is the token-level edit distance between the $t$-th token of $o_{i}$ and $s_{j}$, and $T$ represents the total number of tokens.
By minimizing $d(o_{i},s_{j})$ for each log $o_{i}$, we cluster logs around their nearest representative center in $S$, ensuring that logs within a cluster are as similar as possible according to the defined metric.
@@ 2.2 Template Generation
Template generation aims to differentiate variables from fixed components in logs. It starts with finding the Longest Common Subsequence (LCS), which indicates shared elements across logs. Then, each log is compared with the LCS to pinpoint unique tokens and their locations; for example, comparing ABC with the LCS B reveals distinct tokens at positions A-0 and C-2.
For each unique position, we calculate information entropy inline using $H(X)=-\sum P(x_{i})\log_{2}P(x_{i})$, where $P(x_{i})$ is the probability of the token at that position, and compile a list of these tokens.
To classify a position as a variable, we apply the inline decision: a position is variable if $H>\theta$, otherwise fixed, where $H$ represents information entropy and $\theta$ is a set threshold depending on data characteristics and iterative experiments. When $H$ surpasses $\theta$, we substitute all tokens at that position with the variable marker <><*>< ∗ >.
Additionally, we mark digit-dominated tokens as variables <><*>< ∗ > and use NLP to identify fixed-value tokens (e.g., hash values), enhancing variable identification accuracy. This method effectively identifies variables and fixed parts in logs, adapts to different types of logs, and accurately identifies changing parts while preserving the log structure.
@@ 2.3 Three-S-hop Chain-of-Thought Merging
In advanced log message analysis, after categorizing messages into distinct clusters and extracting unique log templates, we encounter a significant challenge: reconciling log templates that are semantically similar but differ in length. Despite structural variations, these templates often represent identical or closely related events. This challenge primarily arises from the limitations of length-based classification methods, which frequently fail to recognize that templates with equivalent semantics but different expressions may correspond to the same event type.
Figure 3: Three-S-hop Chain-of-Thought Merging Method
Illustrated in Figure 3, We propose Three-S-hop Chain-of-Thought Merging using Large Language Models (LLMs) for semantic parsing. This includes data processing and three dialogue rounds: Structure QA (examines structure and length differences), Semantic QA (probes meanings for semantic equivalences), and Solution QA (decides on merging based on prior analyses). This approach merges templates of different lengths but similar meanings through structural and semantic analysis, enhancing accurate identification.
Dataset Template Avg. Len. Various Lengths Proportion Messages
@@
3 Experiment
3.1 Implement Details
Datasets
Experiments are conducted on the most widely-used benchmark datasets published in LogHub[14]. More details are available in Table 1.
Implementation and Configuration
We implement Lemur based on Python 3.10, Apple M3 chip with 8-core CPU, 10-core GPU 16GB RAM, and macOS Sonoma(14.2.1). In experiments, $D$ is the token set for word-splitting, $k$ denotes the number of the $S$ in each bucket, the Jaccard similarity threshold for within-bucket merging $J_{T}$ and the entropy threshold for token identification $\theta$ are shown in Table 2. We use GPT-4 as the base to implement the three-S-hop chain of thought merging, and Lemur is suitable for other LLMs. Lemur uses these models to conduct inference just by locally loading the weights of these models or utilizing the APIs provided by these LLMs.